OpenStudy (anonymous):

find an equation of a line with the following points: (-7,-4);(-7,5)

OpenStudy (anonymous):

x = -7

OpenStudy (anonymous):

Use two point formula: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

If you try to use dat @waterineyes you are going to have a problem

OpenStudy (anonymous):

Or there is no need as x coordinate of both the points are same.. That means line is parallel to y axis.. So: x = -7 would be the equation..

OpenStudy (anonymous):

Because that's a vertical line and its gradient is actualy undefined according to that formular

OpenStudy (anonymous):

Actually Let me try..

OpenStudy (anonymous):

It is not right but you will get the answer though: \[\frac{y - y_1}{y_2 - y_1} = \frac{x - (-7)}{0} \implies x - (-7) = 0 \implies x = -7\]

OpenStudy (anonymous):

i got the answer you did when i solved it

OpenStudy (anonymous):

you're cheating Maths lol I like it....

OpenStudy (anonymous):

Ha ha ha...

OpenStudy (anonymous):

thanks for the help ^.^/

OpenStudy (anonymous):

Welcome dear..